Monday Operations Meeting : 13/4/2015¶
Agenda:¶
- Action Items from previous meeting
- Document RATS daemon crash signs in beamline documentation (Henry)
- minor changes to checklists (Steve/Henry) : (Stands)
- Chase the OnMon changes in the MDR action items (Steve) : (Stands)
- Work on being able to access archiver data from external sites (Paolo/Pierrick) : (Stands)
- Get input on run plan from Jaroslaw (Jaroslaw/Steve/Milorad) : (Stands)
- Work towards replacement of run conditions spreadsheet (All)
- Review of weekend run
- Actions from previous runs
- Overview of readiness list from MDR1 and MDR2 : http://micewww.pp.rl.ac.uk/projects/operations/wiki/Operations_Readiness_To_Do_List
- Review of web whiteboard : http://micewww.pp.rl.ac.uk/projects/operations/wiki/Web_Whiteboard
- Plans for April 18/19 running
- Data taking plans : http://micewww.pp.rl.ac.uk/projects/operations/wiki/RunPlan20150418
- Detector readiness : TOF2, KL, EMR and DAQ to read them out
- Training plans
- Shifters, BLOCs and SOCs
- Friday training
- Longer term plans
- 18/19Apr15 and 25/26Apr15: DS, CKOV,TOF0,TOF1, EMR, KL, TOF2 (& tracker?)
- Check the TOF calibration. It will be important to take some data to be sure that the calibration is stable;
- Do any other calibration runs required for the PID detectors
- First pass of MICE Muon Beam (MMB) commissioning:
- Start with the currents used in Step I to check the degree of reproducibility of particle rates;
- Based on the revised simulation of the beam line (Jaroslaw Pasternak), start the search for the optimum magnet settings;
- Commissioning and calibration of TOF2, KL and EMR will begin as soon as the south-side PRY is complete and the magnets and downstream detectors are in their final Step IV positions.
- 18/19Apr15 and 25/26Apr15: DS, CKOV,TOF0,TOF1, EMR, KL, TOF2 (& tracker?)
- AOB
- Next meeting : 3pm Monday 20th April
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Minutes of Monday Operations Meeting : 13/4/2015
Present : Milorad, Chris R, Henry, Pierrick, Adam, Jaroslaw, Steve¶
- Action Items from previous meeting
- Document RATS daemon crash signs in beamline documentation (Henry)
- This failure mode has been written up and is part of the standard shifter target training
- minor changes to checklists (Steve/Henry) : (Stands)
- being worked on. Will be done by the end of the week.
- Chase the OnMon changes in the MDR action items (Steve) : (Stands)
- Rhys has got in touch promising to finish these items by the weekend.
- Work on being able to access archiver data from external sites (Paolo/Pierrick) : (Stands)
- Paolo and Pierrick are aware of the problem and will discuss it.
- Get input on run plan from Jaroslaw (Jaroslaw/Steve/Milorad) : (Stands)
- See later in the minutes
- Work towards replacement of run conditions spreadsheet (All)
- A link to the CDBViewer has been put in the Operations documentation (http://micewww.pp.rl.ac.uk/projects/operations/wiki/Operations) and shifter documentation (http://micewww.pp.rl.ac.uk/projects/operations/wiki/MICE_Shifter_Guides). It will also go into the
binder in the CR.
- A link to the CDBViewer has been put in the Operations documentation (http://micewww.pp.rl.ac.uk/projects/operations/wiki/Operations) and shifter documentation (http://micewww.pp.rl.ac.uk/projects/operations/wiki/MICE_Shifter_Guides). It will also go into the
- Document RATS daemon crash signs in beamline documentation (Henry)
- Review of weekend run
- There was no data taking this weekend. A problem with the Run Control was noted during the Friday reference run
however. The shift instructions ask the shifter to start a new Run Control. However this could lead to more than one
instance of the Run Control being active which is something that EPICS finds to be Evil.Action Item Steve to ask Chris to insert a line in the shift documentation asking the shifter to make sure
no other run control instance is operating before starting a new one.Another issue with run control is that there are still odd instances when starting a new run after an old one
can lead to the Run Control packing up. The instruction here is to kill the current instance of run control and start
afresh if this happens.Action Item Steve to ask Chris to insert instructions to this effect in the shifter documentation
- As far as we are aware TOF2, the KL and the EMR should be ready to run on the weekend.
No detector experts were present at the meeting to confirm.
- Henry raised an issue with access. Apparently the shifters need to have swipe access to the target area to put the
target in standby mode. We haven't been asking for this - indeed noone here could remember how to do it.Action Item Milorad to check with Debbie/Rose what the appropriate process is to get swipe access, possibly using the
new ISIS Visitor Database.
- Henry flagged a potential concern with the level of the target laser signals. THey seem to be smaller than they used to
be but we don't know if it is a gradual but constant degradation or a sudden decrease. Ed will look at the laser levels
before the weekend. Henry pointed out that this info is not in the archiver and there isn't an alarm so the priority was
to get this into a state where the shift crew will notice if something happens. Pierrick promises to do this next week,
but it won't be ready for this weekend.
- There was no data taking this weekend. A problem with the Run Control was noted during the Friday reference run
- Actions from previous runs
- No change to the web whiteboard
Action Item Pierrick will update his item in the readiness list.
Action Item Steve to update MDR2-CMP-01. This will be available but in the new version of the framework.
- Run plan for April 18/19
- The run plan depends on whether the DS is operational or not. If it is then a proto-run plan (in priority order)
- Do a reference run with pion beam on Friday
- Do measurements with the muon beam and DS that Jaroslaw is working to define now
- Expand the beam and do a TOF calibration over the entire area of the beam
- If the DS is not available, we can still do the last item - although it will take longer.
- Action Item Jaroslaw to define the run plan settings and give them to Milorad by Wednesday
- Celeste will be acting as BLOC and SOC for the weekend run
- Action Item Steve to track down the PPS training status of weekend shifters (none of whom were at the
meeting) to see if we need to do a PPS tutorial on Friday. - Shifters should keep an eye on the run plan and must attend the Friday meeting
- The run plan depends on whether the DS is operational or not. If it is then a proto-run plan (in priority order)
- AOB
- Milorad would like John G, Mark Tucker and Pierrick to be around to liaise with the Daresbury crew who are arriving
to look at the DS tomorrow - We would like to ask Yordan to be available on Friday afternoon (at least) to check that the DAQ can read out everything.
Action Item Milorad to talk to Yordan about this
- Milorad would like John G, Mark Tucker and Pierrick to be around to liaise with the Daresbury crew who are arriving
Updated by Boyd, Steven over 8 years ago ยท 17 revisions